# VaR at Its Extremes: Impossibilities and Conditions for One-Sided Random Variables

### v1

@@ 1 Introduction
A variety of risk measures have been developed to refine or extend VaR. Among them, the conditional tail expectation – also known as expected shortfall – [Acerbi.2002, Tasche2002] addresses several limitations of VaR, most notably the failure of sub-additivity. Nevertheless, interest in VaR has persisted, partly due to its ability to capture the opposite phenomenon of super-additivity, a feature that does not rely on the integrability requirements imposed by alternative risk measures. In this paper, we examine the extremal behaviors of VaR, focusing on the conditions under which it exhibits sub-additivity or super-additivity across all probability thresholds.
@@
Throughout the remainder of the paper, we assume that each random variable $X_{i}$ has support with lower endpoint at zero, that is,
@@
In the final section, we show how this assumption can be relaxed. In particular, we extend all results to the setting where the lower endpoints $a_{i}$ are arbitrary but finite, and also to the complementary case in which the random variables are instead bounded above.
@@ 2 VaR Sub-additivity
$F_{X_{i,k}}(x_{i})$ $=\begin{cases}\dfrac{\mathbb{P}(X_{i}\leq x_{i},\,S\leq k)}{F_{S}(k)},&x_{i}<k,\\[8.50012pt] 1,&x_{i}\geq k,\end{cases}$
$F_{S_{k}}(s)$ $=\begin{cases}\dfrac{F_{S}(s)}{F_{S}(k)},&s<k,\\[4.25006pt] 1,&s\geq k.\end{cases}$
@@
${\mathrm{VaR}}_{p}[S_{k}]={\mathrm{VaR}}_{pF_{S}(k)}[S]$ $\leq\sum_{i=1}^{n}{\mathrm{VaR}}_{pF_{S}(k)}[X_{i}]$
$\leq\sum_{i=1}^{n}{\mathrm{VaR}}_{p}[X_{i,k}],$
@@ 3 VaR super-additivity
$\Phi(x_{1},\dots,x_{n})\geq\Phi(y_{1},\dots,y_{n})\quad\text{whenever }x_{i}<y_{i}\text{ for all }i\in\{1,\dots,n\},$
then $\Phi$ is SD. Moreover, suppose each $\phi_{i}$ is continuous and
@@
Since $x_{i}<s$ for all $i$, the non-increasing property implies
@@
$\Phi(x_{1},\dots,x_{n})=\sum_{i=1}^{n}\phi_{i}(x_{i}),$
with each $\phi_{i}$ continuous. If every $\phi$ is non-increasing then the sum of non-increasing functions is non-increasing i.e. $\Phi$ is non-increasing.
Conversely, suppose that $\Phi$ is non-increasing, but for contradiction, some coordinate function $\phi_{j}$ is not non-increasing i.e. there exists $x^{*}_{j}<y^{*}_{j}$ such that $\phi_{j}(x^{*}_{j})<\phi_{j}(y^{*}_{j})$ for some $j\in\{1,\dots,n\}$. Set $\epsilon=\phi_{j}(y^{*}_{j})-\phi_{j}(x^{*}_{j})>0$ and by continuity of each $\phi_{i}$, $i\neq j$, there exists $\delta_{i}$ such that $|\phi_{i}(y_{i})-\phi_{i}(x_{i})|<\frac{\displaystyle\epsilon}{\displaystyle n}$ whenever $|y_{i}-x_{i}|<\delta_{i}$. From the continuity domain of each $\phi_{i}$, pick $x^{*}_{i}<y^{*}_{i}$ such that $|y^{*}_{i}-x^{*}_{i}|<\delta_{i}$, then for all those $x^{*}_{i}<y^{*}_{i}$ together with $x^{*}_{j}<y^{*}_{j}$,
$\Phi(y^{*}_{1},\dots,y^{*}_{n})-\Phi(x^{*}_{1},\dots,x^{*}_{n})$ $=\sum_{i=1}^{n}\left(\phi_{i}(y^{*}_{i})-\phi_{i}(x^{*}_{i})\right),$
$=\epsilon+\sum_{i\neq j}\left(\phi_{i}(y^{*}_{i})-\phi_{i}(x^{*}_{i})\right),$
$>\epsilon-\sum_{i\neq j}\frac{\displaystyle\epsilon}{\displaystyle n},$
$=\frac{\displaystyle\epsilon}{\displaystyle n},$
$>0.$
i.e. $\Phi$ is not non-increasing, leading to a contradiction. Hence each $\phi_{i}$ must be non-increasing. Consequently, by the first part of the proof, $\Phi$ is SD whenever all $\phi_{i}$ are non-increasing. ∎
@@
it suffices to verify that each $\phi_{i}(x_{i})$ is non-increasing on $0,\infty)$. This is convenient, as it reduces the verification of SD to checking each margin separately. The following example gives instances of marginal distributions for which the non-increasing property holds.
@@
We present below several standard continuous distribution functions $F_{X_{i}}$ for which the non-increasing property of $\phi_{i}$ holds.
@@
    the classical Mills ratio bound [Mills1926]. Thus $\phi_{i}^{\prime}(x_{i})\leq 0$ for all $x_{i}\geq 0$.
@@
    $F_{X_{i}}(x_{i})=\exp\!\left\{-\frac{\displaystyle\log(1+x_{i})^{\alpha_{i}}}{\displaystyle x_{i}}\right\},\qquad x_{i}\geq 0,\ \alpha_{i}\in(-\infty,1),$
@@
Although the non-increasing property is tractable, it is stronger than what is required for VaR super-additivity. The next example shows that $\Phi$ may be SD without being non-increasing.
@@
$,F_{X_{3}}(x_{3})$ $=\begin{cases}\dfrac{x_{3}^{2}}{e},&0\leq x_{3}\leq 1,\\[5.69054pt] \exp\!\left(-\dfrac{1}{\sqrt{x_{3}}}\right),&x_{3}>1.\end{cases}$

### v2

@@ 1 Introduction
A variety of risk measures have been developed to refine or extend VaR. Among them, the conditional tail expectation – also known as expected shortfall – [Acerbi.2002, Tasche2002] which addresses several limitations of VaR, most notably the failure of sub-additivity. Nevertheless, interest in VaR has persisted, partly due to its ability to capture the opposite phenomenon of super-additivity, a feature that does not rely on the integrability requirements imposed by alternative risk measures. In this paper, we examine the extremal behaviors of VaR, focusing on the conditions under which it exhibits sub-additivity or super-additivity across all probability thresholds.
@@
Throughout Sections 2 and 3 of the paper, we assume that each random variable $X_{i}$ has support with lower endpoint at zero, that is,
@@
In the final section, Section 4, we show how this assumption can be relaxed. In particular, we extend all results to the setting where the lower endpoints $a_{i}$ are arbitrary but finite, and also to the complementary case in which the random variables are instead bounded above.
@@ 2 VaR Sub-additivity
$F_{X_{i,k}}(x_{i})=\begin{cases}\dfrac{\mathbb{P}(X_{i}\leq x_{i},\,S\leq k)}{F_{S}(k)},&x_{i}<k,\\[8.50012pt] 1,&x_{i}\geq k,\end{cases}\quad\mathrm{and}\quad F_{S_{k}}(s)=\begin{cases}\dfrac{F_{S}(s)}{F_{S}(k)},&s<k,\\[4.25006pt] 1,&s\geq k.\end{cases}$
@@
${\mathrm{VaR}}_{p}[S_{k}]={\mathrm{VaR}}_{pF_{S}(k)}[S]\leq\sum_{i=1}^{n}{\mathrm{VaR}}_{pF_{S}(k)}[X_{i}]\leq\sum_{i=1}^{n}{\mathrm{VaR}}_{p}[X_{i,k}],$
@@ 3 VaR super-additivity
$\Phi(x_{1},\dots,x_{n})\geq\Phi(y_{1},\dots,y_{n})\quad\text{whenever }x_{i}\leq y_{i}\text{ for all }i\in\{1,\dots,n\},$
then $\Phi$ is SD. In particular, if $\Phi$ can be written as
@@
Since $x_{i}\leq s$ for all $i$, the non-increasing property implies
@@
$\Phi(x_{1},\dots,x_{n})=\sum_{i=1}^{n}\phi_{i}(x_{i}).$
If every $\phi$ is non-increasing then the sum of non-increasing functions is non-increasing i.e. $\Phi$ is non-increasing.
Conversely, suppose that $\Phi$ is non-increasing. For each $i\in\{1,\dots,n\}$, take $x_{i}\leq y_{i}$ and set $x_{j}=y_{j}=z$, $\forall j\neq i$, then
$\Phi(x_{1},\dots,x_{n})=\phi_{i}(x_{i})+\sum_{j\neq i}\phi_{i}(z)\geq\Phi(y_{1},\dots,y_{n})=\phi_{i}(y_{i})+\sum_{j\neq i}\phi_{i}(z),$
$\implies\phi_{i}(x_{i})\geq\phi_{i}(y_{i}),$
i.e. each $\phi_{i}$ is non-increasing. Consequently, by the first part of the proof, $\Phi$ is SD whenever all $\phi_{i}$ are non-increasing. ∎
@@
it suffices to verify that each $\phi_{i}(x_{i})$ is non-increasing on $0,\infty)$. This is convenient, as it reduces the verification of SD to checking each margin separately. The next example lists several standard continuous marginal distributions $F_{X_{i}}$ for which the function $\phi_{i}$ indeed has the required monotonicity property.
@@
We collect below several familiar continuous distributions whose associated functions $\phi_{i}$ are non-increasing.
@@
    the classical Mills ratio bound [Mills1926]. Thus $\phi_{i}^{\prime}(x_{i})\leq 0$ for all $x_{i}\geq 0$ i.e. $\phi_{i}$ is non-increasing on $0,\infty)$.
@@
    $F_{X_{i}}(x_{i})=\exp\!\left(-\frac{\displaystyle\log(1+x_{i})^{\alpha_{i}}}{\displaystyle x_{i}}\right),\qquad x_{i}\geq 0,\ \alpha_{i}\in(-\infty,1),$
@@
Although the non-increasing property is tractable, it is stronger than what is required for VaR super-additivity. The next example shows that $\Phi$ may be SD without, or equivalently without each $\phi_{i}$, being non-increasing.
@@
$F_{X_{3}}(x_{3})$ $=\begin{cases}\dfrac{x_{3}^{2}}{e},&0\leq x_{3}\leq 1,\\[5.69054pt] \exp\!\left(-\dfrac{1}{\sqrt{x_{3}}}\right),&x_{3}>1.\end{cases}$